OpenStudy (anonymous):

The radius of a sphere is increasing at a rate of 4mm/s. How fast is volume increasing when the diameter is 80mm.

OpenStudy (anonymous):

You start off knowing dr/dt and you want to find out dv/dt

OpenStudy (anonymous):

To get that you can just do \(\Large \frac{dr}{dt}*\frac{dV}{dr}\) Do you know how to get dV/dr?

OpenStudy (anonymous):

dV would just be the derivative of the volume correct?

OpenStudy (anonymous):

Well, tempest, we're dealing with related rates, so it's not enough anymore to say just "the derivative." We have to say "derivative with respect to..." dV/dr means derivative of volume with respect to the radius. dV/dt means derivative of volume with respect to time.

OpenStudy (anonymous):

Ok. I have the solution to the problem and I am confused on how they simplified dV/dt It is number 6

OpenStudy (anonymous):

What they have there is dV/dt = dV/dr * dr/dt

OpenStudy (anonymous):

The diameter is 80, so for the radius, they plugged in r = \(\frac{1}{2}*80\)

OpenStudy (anonymous):

Because the radius is half. Ok I get that

OpenStudy (anonymous):

and \(\Large \frac{4}{3}\pi*3\) became \(\Large 4\pi\) And of course, the problem told you that dr/dt was 4, so they plugged 4 in for dr/dt.

OpenStudy (anonymous):

So basically, I just needed to find the derivitive of the volume with respect to time, and multiply that with the deriv of the radius with respect to time?

OpenStudy (anonymous):

Because the equation of the volume has radius in ti so that is how they are related?

OpenStudy (anonymous):

What you said just now is... not quite right. The problem asks you for the derivative of the volume with respect to time, but it's not so simple to just find that. However, we know that V = (4/3)pi*r^3 so, from that equation, it's easy to find derivative with the volume with respect to RADIUS.

OpenStudy (anonymous):

Once we know derivative of the volume with respect to radius, we can multiply that by dr/dt to get dV/dt like we wanted.

OpenStudy (anonymous):

How do we get the deriviative of dV/dr

OpenStudy (anonymous):

You have the equation V = (4/3) pi*r^3 Take the derivative with respect to r.

OpenStudy (anonymous):

Use exponent rule.

OpenStudy (anonymous):

v' = 3r^2? I am unsure about the (4/3)pi

OpenStudy (anonymous):

That just gets multiplied by the 3.

OpenStudy (anonymous):

derivative of ax^n is an*x^(n-1)

OpenStudy (anonymous):

So no need to take the deriviative of 4/3 pi because it will get multiplied by 3

OpenStudy (anonymous):

That part of the equation is just a constant.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

So dV/dr = (4/3)pi*3r^2 = 4pi*r^2

OpenStudy (anonymous):

That is what we are solving for right? I think i am making this harder then it needs to be. I have been stumped on this problem for a good 20 minutes...

OpenStudy (anonymous):

lol well perhaps I'm not explaining it super simply.

OpenStudy (anonymous):

The radius of a sphere is increasing at a rate of 4mm/s. How fast is volume increasing when the diameter is 80mm. "Radius of a sphere is increasing at a rate..." dr/dt = 4 "How fast is volume increasing?" We need to find dV/dt \(\Large \frac{dr}{dt}*\frac{dV}{dr} = \frac{dV}{dt}\) So we have a way to find it, but we must first find dV/dr. How do we find dV/dr? V=(4/3)pi*r^3 Derive with respect to r.

OpenStudy (anonymous):

So since dr/dt = 4 and dV/dt = 4pi*r^2, I just plug in 4 into r?

OpenStudy (anonymous):

Nonono. dV/dr = 4pi*r^2 You haven't found dV/dt yet.

OpenStudy (anonymous):

Ok so how do we find it?

OpenStudy (anonymous):

\(\Large \frac{dr}{dt}*\frac{dV}{dr} = \frac{dV}{dt}\)

OpenStudy (anonymous):

oh... dr/dt is 4 and dV/dr is 4pi r^2. We are solving for when the dimaeter is 80. So r is = 40 and then * dV/dr by dr/dt

OpenStudy (anonymous):

when dr/dt is 4

OpenStudy (anonymous):

Right, I believe.

OpenStudy (anonymous):

Ok i got it correct.

OpenStudy (anonymous):

So basiclly, find a formula that related both derivitves, and derive that formula respect to one of the varibles. ex: dx/dt and dy/dt. You need to relate using a formula derivitive dy/dx

OpenStudy (anonymous):

Can someone confirm?